West Fraser Timber Company (WFT) is expected to have free cash flow in the coming year of $2 million and its free cash flow is expected maintain at a sustainable growth rate of 4% per year. It has a debt worth $10 million. It’s equity cost of capital is 12%, cost of debt before tax is 6%, and it pays a corporate tax rate of 30%. If WFT Company maintains a debt-equity ratio of 0.5 and the company has 3 million common shares outstanding, what is the fair value of WFT stock?
Answer #1

Solution :

Free cash flow = $2 million

Growth rate = 4%, Debt = $10 million

Cost of debt = 6% or 0.06

Equity cost of capital = 12% or 0.12

Tax rate = 30% or 0.30

Working 1: Calculation of equity

Debt equity ratio = 0.5

Debt / Equity = 0.5

$10 million / Equity = 0.5

Equity = $10 million / 0.5

Equity = $20 million

Working 2: Weighted average cost of capital (WACC)

WACC = (Debt / (Equity + Debt) * Cost of debt * (1 - tax rate)) + (Equity / (Equity + Debt) * Cost of equity))

Using given details ,

WACC = ($10 million/($20 million + $10 million) * 0.06 * (1 - 0.30)) + ($20 million / $20 million + $10 million) * 0.12

WACC = (0.3333 * 0.06 * 0.70) + (0.6667 * 0.12)

WACC = 0.0140 + 0.080 = 0.094 or 9.4%

Working 3: Calculation of value of WFT

Value of WFT = Expected free cash flows / (WACC - Growth rate

Here,

Expected free cash flows = $2 million

WACC (as per working 2) = 0.094

Growth rate = 4% or 0.04

Now put the values into the above formula:

Value of WFT = $2 million / (0.094 - 0.04)

Value of WFT = $37.04 million

Working 4: Calculation of fair value of WFT stock

Fair value of stock = Value of WFT / Outstanding common shares

Here,

Value of WFT = $37.04 million

Outstanding common shares = 3 million shares

Now,

Fair value of stock =$37.04 million/3 million shares

Fair value of stock = $12.35 per share

Note: Figures are rounded off upto 4 decimals.